# NP-Hardness and ETH-Based Inapproximability of Communication Complexity via Relaxed Interlacing

### v1

@@ Abstract
Only constant alternations required. The optimal protocol on the instances our reduction maps onto require only constant round alternation. This proves NP-hardness for unrestricted rounds of alternation on a more restrictive class of functions than those with optimal protocols requiring a growing number of rounds of communication. This also separates our result from recent concurrent and independent work by Hirahara et al. [3], whose hardness proof relies on unbounded alternations.
@@
footnotetext: Concurrent and independent work by Hirahara, Ilango, and Loff [3] also establishes NP-hardness of deciding $D(f)\geq k$, but their reduction requires an unbounded number of alternations and does not extend to ETH-based additive inapproximability.
@@ 1.1 Notation & Parameters
$\varepsilon_{q,t}$ AGHP error of (q,t)(q,t)‑balanced set $\varepsilon_{q,t}:=(qt|C|)^{-c}$ §2 Drives $y^{\prime}$ in Lem. 10
$q$ total outer blocks $q\geq t$; $q=rt$ in Lem. 14 Def. 10
$t$ # of fully independent blocks $1\leq t\leq 2^{b}$; power of two; $\Delta\leq\log t$ Lem. 10 Lem. 13 Inner interlace factor
$r$ replication with $q=rt$ $r\geq 1$ (int.) Lem. 14 Adds $\lceil\log r\rceil$ to budget $B$
@@
$B^{\prime}$ One round budget in Lem. 14 $B=\log t+\lceil\log r\rceil$ Lem. 14 Any optimal protocol must use all $B$ bits
@@ 2.1 Almost k-Wise Independent Sets
    an integer $q\geq 1$ (number of interlacing layers),
@@
    a balance threshold $t\geq 1$ (number of columns we control),
@@
||= $\mathcal{B}|\;=\;2^{\,O\!\left(\log\log(q\ell)+k+\log(1/\varepsilon_{q,t})\right)}\;=\;(qp)^{\,O(t)}\cdot\operatorname{polylog}(q),$ (3)
such that the uniform distribution on $\mathcal{B}$ is $\varepsilon_{q,t}$-close (in statistical distance) to $k$-wise independence. Importantly, this set can be constructed in $(qp)^{\,O(t)}\cdot\operatorname{polylog}(q)$ time.
@@
The choice $\varepsilon_{q,t}:=(qt|C|)^{-c}$ is far smaller than needed in every later bound, yet keeps $|\mathcal{S}|$ polynomial in all input parameters. Hence we dispense with a user-supplied accuracy parameter for the rest of the paper.
@@ 2.2 The Relaxed Interlace Operation
$|[q]\times R|\cdot|\mathcal{S}|\;=\;q\,|R|\,\cdot\,2^{\,O\!\left(\log\log(q\ell)+t\ell+\log(1/\varepsilon)\right)}$
@@ 2.3 Structural Lemmas on Relaxed Interlace
$y^{\prime}\;:=\;1-\left(1+\varepsilon\right)\,(1-y)\;=\;y-\varepsilon(1-y)\;=\;(1+\varepsilon)\,y-\varepsilon.$
@@
Step 1: trimming the rows. Because $R^{\prime}$ is (Q,T)(Q,T)‑equipartitioned and $|Q|=t$, choose exactly $T$ rows from each block $\{q\}\times R$ with $q\in Q$ and discard every other row. Denote the surviving set by $R^{\ast}\subseteq R^{\prime}.$ Then $R^{\ast}$ is an exact (Q,T)(Q,T)‑equipartition, hence $\left|R^{\ast}\right|=tT$ and condition (i) of Definition 8 with parameter $x$ is met.
Step 2: projecting the columns. Let $\pi_{Q}:\mathcal{S}\to C^{t}$ be the natural projection that keeps the $t$ coordinates indexed by $Q$. Write $D:=\pi_{Q}(C^{\prime})\subseteq C^{t}.$ Because $C^{\prime}$ is a sub‑multiset of $\mathcal{S}$, $D$ inherits multiplicities. By the balance Lemma 8, the uniform distribution on $\mathcal{S}$ is $\varepsilon$‑close to tt‑wise independence; equivalently, every pattern in $C^{t}$ occurs in $\mathcal{S}$ with frequency $\frac{1}{|C|^{\,t}}\pm\varepsilon$. A standard averaging argument shows that at most an $\varepsilon$‑fraction of the tuples in $C^{t}$ are undersupplied compared with the ideal uniform frequency. Keeping only the well‑supplied patterns,
$D^{\ast}\;:=\;\left\{\,\mathbf{c}\in C^{t}\;\big{|}\;\text{frequency of $\mathbf{c}$ in $C^{\prime}$ }\geq\left(1-\varepsilon\right)\frac{|C^{\prime}|}{|C^{t}|}\right\},$
we lose at most an $\varepsilon$‑fraction of columns from $C^{\prime}$. More precisely,
$|D^{\ast}|\;\geq\;|C^{t}|\;\left(y-\varepsilon(1-y)\right)\;=\;|C^{t}|\,y^{\prime}.$
Step 3: lifting back to $\mathcal{S}$. For every pattern $\mathbf{c}\in D^{\ast}$ keep one representative column in $C^{\prime}$ whose projection is $\mathbf{c}$; call the resulting set $C^{\ast}\subseteq C^{\prime}$. By construction $|C^{\ast}|=|D^{\ast}|\geq|C^{t}|\,y^{\prime}$.
Step 4: identifying the subgame. Re‑index the rows in $R^{\ast}$ as $([t]\times R)\ni(i,r)$ with $i$ the coordinate in $Q$ (keeping the original order), and re‑index each column in $C^{\ast}$ by its QQ‑projection in $C^{t}$. Under that relabelling the submatrix $N^{\ast}:=\varepsilon\!\left(N;R^{\ast},C^{\ast}\right)$ is exactly the classical interlace $M\langle t\rangle$ restricted to $R^{\ast}_{\text{canon}}\times D^{\ast}.$ The row set is an exact TT‑equipartition, and $|D^{\ast}|=\lceil|C|^{t}\,y^{\prime}\rceil$ by the previous step, hence $N^{\ast}\in M\left[t,x,y^{\prime}\right].$
@@ 3.1 Lower Bounding the Complexity
$(1+\varepsilon)h^{\prime}(t)\cdot 2^{-B}r-\varepsilon$ $\geq h(t)$ (7)
$D\left(M\left[1,2^{-B},\left((1+\varepsilon)\,h^{\prime}(t)\,2^{-B}-\varepsilon\right)^{1/t}\right]\right)\;$  $\geq\;1,$ (8)
@@
    There exists $i\in[B]$ for which $P(i)$ is true and $P(i+1)$ is false.
@@ 2. Step 2:
    $N^{\ast}\;\in\;M\left[t,2^{-B},y^{\prime}\right],\qquad y^{\prime}\;=\;(1+\varepsilon)\,h^{\prime}(t)\,2^{-B}-\varepsilon.$
@@ 3. Step 3:
    $D\left(N^{\ast}\right)\;\geq\;D\left(M\left[1,2^{-B},\left((1+\varepsilon)\,h^{\prime}(t)\,2^{-B}-\varepsilon\right)^{1/t}\right]\right).$
@@ 4. Step 4:
    $D\left(M\left[1,2^{-B},\left((1+\varepsilon)\,h^{\prime}(t)\,2^{-B}-\varepsilon\right)^{1/t}\right]\right)\;\geq\;1.$
@@
Case b. The row–block invariant is broken at the first step i+1i{+}1 ($1\leq i<B$).
@@ 1. Step 1:
    To break the invariant at step i+1i{+}1 the protocol must have asked at least $\left\lceil\log r\right\rceil$ row questions (each halves the number of viable blocks). Hence no more than $B-\left\lceil\log r\right\rceil$ bits have attacked the columns, so
    $\frac{|C_{i}|}{|\mathsf{Cols}(\widehat{M})|}\;\geq\;h^{\prime}(t)\;2^{-(B-\lceil\log r\rceil)}\;\geq\;h^{\prime}(t)\,r\,2^{-B}\;=:\;y.$
@@ 2. Step 2:
    $t=|Q|,\qquad x=2^{-b},\qquad y=h^{\prime}(t)\,r\,2^{-B}.$
@@
    $N^{\ast}\;\in\;M\left[t,2^{-b},y^{\prime}\right],\qquad y^{\prime}=(1+\varepsilon)\,h^{\prime}(t)\,r\,2^{-B}-\varepsilon.$
@@ 3.2 Near-Exact Separation of Interlaced Functions
Lemma 14.
Fix integers $b\geq 1$, $r\geq 1$ and let $t\leq 2^{b}$ be a power of two. Choose $0<\delta\leq\tfrac{1}{\sqrt{2}}-\tfrac{1}{2}.$
Assume $M$ is $(\delta,b)$‑robust and that
$D\left(M\left[2,2^{-\log{r}-\log{t}},\left((1+\varepsilon)\,2^{-3\log{r}-3\log{t}}-\varepsilon\right)^{1/t}\right]\right)\;$ $\geq\;D\left(M\right)+1,$ (9)
@@
$\dfrac{|C^{\prime}|}{|\mathsf{Cols}(\widehat{M})|}\geq\tfrac{1}{2}+\delta.$ (10)
Put $N:=\varepsilon\left(\widehat{M};\mathsf{Rows}(M),C^{\prime}\right)$ and $B^{\prime}:=\log t+\lceil\log r\rceil$.
For every integer $c\geq 1$, every deterministic protocol on $N$ that sends at most $B^{\prime}$ bits but spends fewer than $B^{\prime}-c$ of them on the outer block index leaves a transcript leaf rectangle $R_{c}\times C_{c}$ such that
$D\left(\varepsilon\left(\widehat{M};R_{c},C_{c}\right)\right)\;\geq\;D\left(M\right)+1.$
Hence any protocol attaining the tight bound $D\left(M\right)+\log t+\lceil\log r\rceil$ must devote all $B^{\prime}$ block bits to isolating the $q=rt$ interlaced games.
@@
Let $B^{\prime}=\log t+\lceil\log r\rceil$ and fix an integer $1\leq c\leq B^{\prime}$. Write $s:=B^{\prime}-c$ for the number of questions the protocol does ask about the outer block index.
Step 1: shape of the rectangle after $s$ block bits.
@@
Row side. Since $q=rt$ we can lower bound it by $2^{B^{\prime}-1}+1$. Since $B^{\prime}-c$ bits are spent separating the outer indices, at least one element of the partition must contain $\left\lceil\frac{2^{B^{\prime}-1}+1}{2^{B^{\prime}-c}}\right\rceil$ outer indices with at least $2^{-B^{\prime}}|\mathsf{Rows}(M)|$ inner indices present. Since $c\leq B^{\prime}$ this simplifies to $2^{c-1}+1$ outer indices that must be contained. This means that we can find a $Q,T$-equipartition $R_{c}$ with $|Q|\geq 2^{c-1}+1$ and $T\geq 2^{-B^{\prime}}$.
@@
$|C_{c}|\;\geq\;\frac{|C^{\prime}|}{2^{c}}\;\geq\;\frac{(\tfrac{1}{2}+\delta)\,|\mathsf{Cols}(\widehat{M})|}{2^{\,c}},$ (11)
@@
Extraction. If the protocol does not separate all outer ranks, we obtain rectangle $R_{c}\times C_{c}$. Denoting $g(c,t):=\min{\left(2^{c}-1+1,t\right)}$, we may use Lemma 10 to infer
$N_{c}:=\varepsilon\left(\widehat{M};R_{c},C_{c}\right)\;\in\;M\left[g(c,t),\,2^{-B^{\prime}}\,,\displaystyle y_{c}:=\frac{\frac{1}{2}+\delta}{2^{\,c}}\right].$ (12)
Step 2: project onto two blocks (Condition 5).
Apply the (Extended Maximum Projection Lemma). with $l=2$:
$N_{c}^{\prime}\in M\left[2,2^{-\log{r}-\log{t}},\left(y_{c}\right)^{\,\frac{2}{{g(c,t)}}}\right].$
$\left(\dfrac{\frac{1}{2}+\delta}{2^{c}}\right)^{\frac{2}{g(c,t)}}$ $=\left(\frac{1}{2}+\delta\right)^{\frac{2}{g(c,t)}}\cdot 2^{-\frac{2c}{g(c,t)}}$
Given that $\frac{2}{g(c,t)}$ tends to zero the first term in the product goes to $1$ whilst
$2^{\frac{-2c}{g(c,t)}}\geq\left(2^{-2\log{r}-2\log{t}}\right)^{\frac{1}{t}}\geq\left((1+\varepsilon)\,2^{-3\log{r}-3\log{t}}-\varepsilon\right)^{\frac{1}{t}}.$
Therefore, condition 9 implies
$D\left(N_{c}^{\prime}\right)\;\geq\;D\left(M\right)+1.$
Step 3: monotonicity back to the protocol leaf. Because $N_{c}^{\prime}$ is a submatrix of the leaf rectangle $\widetilde{R}\times\widetilde{C}$, the Monotonicity Lemma 1 gives
$D\left(\varepsilon\left(\widehat{M};R_{c},C_{c}\right)\right)\;\geq\;D\left(N_{c}^{\prime}\right)\;\geq\;D\left(M\right)+1.$
This gives our gap. ∎
Theorem 2 (Near-Exact Separation Theorem).
Given a $(\delta,b)$-robust matrix $M$ such that
$D\left(M\left[2,2^{-\log{r}-\log{t}},\left((1+\varepsilon)\,2^{-\log{r}-\log{t}}-\varepsilon\right)^{1/t}\right]\right)\;$ $\geq\;D\left(M\right)+1,$ (13)
[…]
$N\in\varepsilon\left(\widehat{M};R,C\right).$
Then after $rt$ bits the elements of the partition must all be of the form $R^{\prime}\times\mathsf{Cols}(N)$ where $R^{\prime}\subseteq\mathsf{Rows}(N)$ is a $(Q,|\mathsf{Rows}(M)|(1-2^{-b}rt))$-equipartition for some $Q\subseteq[rt]$ and $|Q|=1$.
Proof.
Using Lemma 14, we may assume that the number of ”wasted bits” $c$ is 0, and hence no bit is spent partitioning the columns. From Condition 5, it we cannot have an element with $2^{-b}|\mathsf{Rows}(N)|$ rows from $2$ different outer indices. Since the number of outer indices equals the number of protocol leaves after rtr\,t bits, for any given outer index $i$ there is exactly one leaf $L_{i}=R_{i}\times C$ containing more than $|\mathsf{Rows}(N)|2^{-b}$ of the form $(i,e)$ with $e\in Rows(M)$. Since all rows must be mapped to a leaf, and there are rtr\,t leaves, this implies that for all $i$, $L_{i}=R_{i}\times C$ where $R_{i}$ contains at least $|\mathsf{Rows}(M)|\left(1-2^{-b}rt\right)$ rows of the form $(i,e)$ with $e\in\mathsf{Rows}(M)$. ∎

### v2

@@ Abstract
footnotetext: Concurrent and independent work by Hirahara, Ilango, and Loff [2] also establishes NP-hardness of deciding $D(f)\geq k$, but their reduction does not address additive inapproximability.
@@ 1.1 Notation & Parameters
$\varepsilon_{q,t}$ AGHP error of (q,t)(q,t)‑balanced set $\varepsilon_{q,t}:=(qt)^{-c}$ §2 Drives $y^{\prime}$ in Lem. 10
$q$ total outer blocks $q\geq t$; $q=rt$ in Thm. 2 Def. 10
$t$ # of fully independent blocks $1\leq t\leq 2^{b}$; power of two; $\Delta\leq\log t$ Lem. 10,13 Inner interlace factor
$r$ replication with $q=rt$ $r\geq 1$ (int.) Thm. 2 Adds $\lceil\log r\rceil$ to budget $B$
@@
$B^{\prime}$ One round budget in Thm. 2 $B^{\prime}=\log t+\lceil\log r\rceil$ Thm. 2 Any optimal protocol must use all $B$ bits
@@ 2.1 Almost $k$-wise Independent Sets
    an integer $q\geq 1$, (number of interlacing layers)
@@
    a balance threshold $t\geq 1$, (number of columns we control)
@@
||= $\mathcal{B}|\;=\;2^{\,O\!\left(\log\log(q\ell)+k+\log(1/\varepsilon_{q,t}p^{t})\right)}\;=\;O(\log(q\ell))\cdot\ell^{\,O(t)}\cdot\operatorname{poly}(q\cdot p^{t}),$ (3)
such that the uniform distribution on $\mathcal{B}$ is $\varepsilon_{q,t}$-close (in statistical distance) to $k$-wise independence. Importantly, this set can be constructed in $O(|\mathcal{B}|)$ time.
@@
The choice $\varepsilon_{q,t}:=(qt)^{-c}$ is far smaller than needed in every later bound, yet keeps $|\mathcal{S}|$ polynomial in all input parameters. Hence we dispense with a user-supplied accuracy parameter for the rest of the paper.
@@ 2.2 The Relaxed Interlace Operation
$|[q]\times R|\cdot|\mathcal{S}|\;=\;q\,|R|\,\cdot\,2^{\,O\!\left(\log\log(q\ell)+t\ell+\log(1/\varepsilon p^{t})\right)}$
@@ 2.3 Structural Lemmas on Relaxed Interlace
$y^{\prime}$ $=\frac{y}{1+\varepsilon}.$
@@
Step 1: trimming the rows. Because $R^{\prime}$ is (Q,T)(Q,T)‑equipartitioned and $|Q|=t$, choose exactly $T$ rows from each block $\{q\}\times R$ with $q\in Q$ and discard every other row. Denote the surviving set by $R^{\ast}\subseteq R^{\prime}.$ Then $R^{\ast}$ is an exact (Q,T)(Q,T)‑equipartition, hence $\left|R^{\ast}\right|=tT$ and condition (i) of 8 with parameter $x$ is met.
Step 2: projecting the columns. Let $\pi_{Q}:\mathcal{S}\to C^{t}$ be the natural projection that keeps the $t$ coordinates indexed by $Q$. Write $D:=\pi_{Q}(C^{\prime})\subseteq C^{t}.$ Because $C^{\prime}$ is a sub‑multiset of $\mathcal{S}$, $D$ inherits multiplicities, by the balance Lemma 8, the uniform distribution on $\mathcal{S}$ is $\varepsilon$‑close to tt‑wise independence; equivalently, every pattern in $C^{t}$ occurs in $D$ with frequency $\frac{1}{|C|^{\,t}}\pm\varepsilon=\frac{1\pm\varepsilon}{|C|^{\,t}}$. Each pattern occurs in $\pi_{Q}(\mathcal{S})$ with frequency at most $\frac{1+\varepsilon}{|C|^{\,t}}$, $D$ contains at least contains $\frac{|C|^{t}}{1+\varepsilon}y$ unique patterns.
More precisely, let $C^{\ast}$ be the set that contains one representative of each pattern in $D$; then
$\frac{|C^{\ast}|}{|C|^{t}}\geq\frac{y}{1+\varepsilon}=:y^{\prime}.$
Step 3: identifying the subgame. Re‑index the rows in $R^{\ast}$ as $([t]\times R)\ni(i,r)$ with $i$ the coordinate in $Q$ (keeping the original order), and re‑index each column in $C^{\ast}$ by its QQ‑projection in $C^{t}$. Under that relabelling the submatrix $N^{\ast}:=\varepsilon\!\left(N;R^{\ast},C^{\ast}\right)$ is exactly the classical interlace $M\langle t\rangle$ restricted to $R^{\ast}\times C^{\ast}.$ The row set is an exact TT‑equipartition, hence $N^{\ast}\in M\left[t,x,y^{\prime}\right].$
@@ 3.1 Lower Bounding the Complexity
$h^{\prime}(t)2^{-B}r/(1+\varepsilon)$ $\geq h(t)$ (7)
$D\left(M\left[1,2^{-B},\left(h^{\prime}(t)\,2^{-B}/(1+\varepsilon)\right)^{1/t}\right]\right)\;$  $\geq\;1,$ (8)
as well as
$\forall k,c,\;D\left(M\left[1,2^{-B},\left(\frac{h^{\prime}(t)2^{-c}}{1+\varepsilon}\right)^{\frac{1}{t}}\right]\right)\geq D\left(M\right)-k$ (9)
with $k\in[B-B^{\prime}]$ and $c\leq\log{t}+k$.
@@
    There exists $i\in[B^{\prime}-1]$ for which $P(i)$ is true and $P(i+1)$ is false.
  3. c)
    There exists $i\in[B^{\prime},B-1]$ for which $P(i)$ is true and $P(i+1)$ is false.
@@ 2. Step 2:
    $N^{\ast}\;\in\;M\left[t,2^{-B},y^{\prime}\right],\qquad y^{\prime}\;=\;h^{\prime}(t)\,2^{-B}/(1+\varepsilon).$
@@ 3. Step 3:
    $D\left(N^{\ast}\right)\;\geq\;D\left(M\left[1,2^{-B},\left(h^{\prime}(t)\,2^{-B}/(1+\varepsilon)\right)^{1/t}\right]\right).$
@@ 4. Step 4:
    $D\left(M\left[1,2^{-B},\left(h^{\prime}(t)\,2^{-B}/(1+\varepsilon)\right)^{1/t}\right]\right)\;\geq\;1.$
@@
Case b. The row–block invariant is broken at the first step i+1i{+}1 ($1\leq i<B^{\prime}$).
@@ 1. Step 1:
    To break the invariant at step i+1i{+}1 the protocol must have asked at least $\left\lceil\log r\right\rceil$ row questions (each halves the number of viable blocks). Hence no more than $B^{\prime}-\left\lceil\log r\right\rceil$ bits have attacked the columns, so
    $\frac{|C_{i}|}{|\mathsf{Cols}(\widehat{M})|}\;\geq\;h^{\prime}(t)\;2^{-(B^{\prime}-\lceil\log r\rceil)}\;\geq\;h^{\prime}(t)\,r\,2^{-B^{\prime}}\;=:\;y.$
@@ 2. Step 2:
    $t=|Q|,\qquad x=2^{-b},\qquad y=h^{\prime}(t)\,r\,2^{-B^{\prime}}.$
@@
    $N^{\ast}\;\in\;M\left[t,2^{-b},y^{\prime}\right],\qquad y^{\prime}=h^{\prime}(t)\,r\,2^{-B^{\prime}}/(1+\varepsilon).$
@@ 4. Step 4:
Case c.
The row–block invariant is broken at the first step i+1i{+}1 ($B^{\prime}\leq i<B$). Let $\widetilde{M}_{i}=\varepsilon\left(\widehat{M};R_{i},C_{i}\right)$ be the unique rectangle that satisfies the invariant immediately before the (i+1)(i{+}1)-st bit is sent.
  1. Step 1:
    Row and column parameters of $\widetilde{M}_{i}$.
    By the invariant $R_{i}$ is a $(Q,T_{0})$-equipartition with
    $t=|Q|,\qquad T_{0}=\left\lceil|R|\,2^{-b}\right\rceil.$
    To break the invariant at step i+1i{+}1 the protocol must have asked at least $\left\lceil\log r\right\rceil$ row questions (each halves the number of viable blocks). Hence no more than $c\leq B-\left\lceil\log r\right\rceil$ bits have attacked the columns, so
    $\frac{|C_{i}|}{|\mathsf{Cols}(\widehat{M})|}\;\geq\;h^{\prime}(t)\;2^{-c}$
  2. Step 2:
    Relaxed $\rightarrow$ classical interlace (Lemma 10). Apply the lemma to $\widetilde{M}_{i}$ with
    $t=|Q|,\qquad x=2^{-B},\qquad y=h^{\prime}(t)\,2^{-c}.$
    $N^{\ast}\;\in\;M\left[t,2^{-B},y^{\prime}\right],\qquad y^{\prime}=h^{\prime}(t)\,2^{-c}/(1+\varepsilon).$
  3. Step 3:
    Apply condition 9.
    Invoking condition 9 with $k=i-B^{\prime}$ gives
    $D\left(N^{\ast}\right)$ $\geq D\left(M\left[t,2^{-B},h^{\prime}(t)2^{-c}\right]\right)$
    $\geq D\left(M\left[1,2^{-B},\left(h^{\prime}(t)2^{-c}\right)^{\frac{1}{t}}\right]\right)$
    $\geq D\left(M\right)-k.$
Implying that case c leads to complexity greater or equalt to the budget. ∎
@@ 3.2 Near-Exact Separation of Interlaced Functions
Theorem 2 (Near-Exact Separation Theorem).
Fix integers $b\geq 1$, $r\geq 1$ and let $t\leq 2^{b}$ be a power of two. Choose $0<\delta\leq\tfrac{1}{\sqrt{2}}-\tfrac{1}{2}.$ For any
$\delta^{\ast}\geq 4\left(\tfrac{1}{2}+\delta\right)^{3}-\tfrac{1}{2},$ (10)
assume $M$ is $(\delta,b)$‑robust and let $B^{\prime}=\log t+\left\lceil\log r\right\rceil$, and $B=B^{\prime}+D\left(M\right)$, such that
$D\left(M\left[2,2^{-b},\left(2^{-B^{\prime}}/(1+\varepsilon)\right)^{2/t}\right]\right)\;$  $\geq\;1,$ (11)
as well as
$\forall k,c,\;D\left(M\left[2,2^{-B},\left(\frac{2^{-c}}{1+\varepsilon}\right)^{\frac{2}{t}}\right]\right)\geq D\left(M\right)-k$ (12)
with $k\in[B-B^{\prime}]$ and $c\leq\log{t}+k$.
@@
$\dfrac{|C^{\prime}|}{|\mathsf{Cols}(\widehat{M})|}\geq\tfrac{1}{2}+\delta^{\ast}.$ (13)
Let $N:=\varepsilon\left(\widehat{M};\mathsf{Rows}(M),C^{\prime}\right)$, and $c$ be an integer such that every deterministic protocol on $N$ that sends at most $B^{\prime}$ bits but spends $B^{\prime}-c$ of them on the outer block index—keep $1-2^{-b}rt$ fraction of rows of each block to the same subgame—leaves a transcript leaf rectangle $R_{c}\times C_{c}$, such that for $c\geq 1$ we have
$D\left(\varepsilon\left(\widehat{M};R_{c},C_{c}\right)\right)\;\geq\;D\left(M\right)+1,$
and for $c=0$, the protocol must isolate all outer indices nearly exactly, we have that $R_{0}$ is a $(Q,|\mathsf{Rows}(N)|(1-2^{-b}rt))$-equipartition for some $Q\subseteq[rt]$ and $|Q|=1$.
@@
Let $B^{\prime}=\log t+\lceil\log r\rceil$ and suppose $1\leq c\leq B^{\prime}$. Write $s:=B^{\prime}-c$ for the number of questions the protocol does ask about the outer block index.
We may then distinguish two cases:
  1. a)
    $\Pi$ exhausts its budget $B^{\prime}$ without ever breaking the invariant.
  2. b)
    There exists $i\in[B^{\prime}-1]$ for which $P(i)$ is true and $P(i+1)$ is false.
  3. c)
    There exists $i\in[B^{\prime},B-1]$ for which $P(i)$ is true and $P(i+1)$ is false.
Let $\widetilde{M}=\varepsilon\left(\widehat{M};R_{B^{\prime}},C_{B^{\prime}}\right)$ be a rectangle that survives after the last communication bit.
  1. Step 1:
    Parameters of the surviving rectangle.
    $R_{B^{\prime}}\text{ is a }(Q,T_{0})\text{-equipartition},\qquad|Q|=t,\qquad T_{0}=\left\lceil|\mathsf{Rows}(M)|\,2^{-b}\right\rceil.$
    Because at most $B^{\prime}$ column bits were revealed,
    $\frac{|C_{B^{\prime}}|}{|\mathsf{Cols}(\widehat{M})|}\;\geq\;2^{-B^{\prime}}$
  2. Step 2:
[…]
    Apply the lemma with $t=|Q|$, $x=2^{-b}$, and the above $y$. It returns subsets $R^{\ast}\subseteq R_{B^{\prime}}$, $C^{\ast}\subseteq C_{B^{\prime}}$ and a subgame
    $N^{\ast}\;\in\;M\left[t,2^{-b},2^{-B^{\prime}}/(1+\varepsilon)\right],$
  3. Step 3:
    Projection to 2 blocks.
    By Lemma 2,
    $D\left(N^{\ast}\right)\;\geq\;D\left(M\left[2,2^{-b},\left(2^{-B^{\prime}}/(1+\varepsilon)\right)^{2/t}\right]\right).$
    According to Equation 11, we have $D\left(N^{\ast}\right)\geq 1$.
Case b. The row–block invariant is broken at the first step i+1i{+}1 ($1\leq i<B^{\prime}$).
Since the invariant is broken, there are at least $\left\lceil\log r\right\rceil$ bits spent on the outer block index, $1\leq c\leq\log t$.
  1. Step 1:
    Shape of the rectangle after $s$ block bits.
@@ 1. Step 1:
    Row side. Since $q=rt$ we can lower bound it by $2^{B^{\prime}-1}+1$. Since $B^{\prime}-c$ bits are spent separating the outer indices, each bit failed to separate outer indices would introduce at least one more interlace for one subgame, so at least one element of the partition must contain $\left\lceil\frac{2^{B^{\prime}-1}+1}{2^{B^{\prime}}-c}\right\rceil$ outer indices with at least $2^{-B^{\prime}}|\mathsf{Rows}(M)|$ inner indices present. Since $c\leq B^{\prime}$ this simplifies to $2^{c-1}+1$ outer indices that must be contained. This means that we can find a $Q,T$-equipartition $R_{c}$ with $|Q|\geq 2^{c-1}+1$ and $T\geq 2^{-B^{\prime}}$.
@@
    $|C_{c}|\;\geq\;\frac{|C^{\prime}|}{2^{c}}\;\geq\;\frac{(\tfrac{1}{2}+\delta^{\ast})\,|\mathsf{Cols}(\widehat{M})|}{2^{\,c}},$ (14)
@@
    Extraction. If the protocol does not separate all outer ranks, we obtain rectangle $R_{c}\times C_{c}$. We may use Lemma 10 to infer
    $N_{c}:=\varepsilon\left(\widehat{M};R_{c},C_{c}\right)\;\in\;M\left[2^{c-1}+1,\,2^{-b-2B^{\prime}}\,,\displaystyle y_{c}:=\frac{\frac{1}{2}+\delta^{\ast}}{2^{\,c}(1+\varepsilon)}\right].$ (15)
  2. Step 2:
    Project onto two blocks (Condition 5). Apply the (Extended Maximum Projection Lemma). with $l=2$:
    $D\left(N_{c}\right)\geq D\left(M\left[2,2^{-B^{\prime}},\left(y_{c}\right)^{\,\frac{2}{2^{c-1}+1}}\right]\right).$
    Let $x=\tfrac{1}{2}+\delta\in\big{[}\tfrac{1}{2},\tfrac{1}{\sqrt{2}}\big{]}$. For any integer $c\geq 1$, since $t\mapsto t^{\frac{2}{2^{c-1}+1}}$ is increasing on $(0,\infty)$,
    $\left(\frac{\tfrac{1}{2}+\delta^{\ast}}{2^{c}}\right)^{\frac{2}{2^{c-1}+1}}\geq\left(\tfrac{1}{2}+\delta\right)^{2}\iff\frac{4x^{3}}{2^{c}}\geq x^{2^{c-1}+1}\iff 4x^{3}\geq 2^{c}x^{2^{c-1}+1}.$
    It remains to verify the last inequality. Set $g_{c}(x):=2^{c}x^{2^{c-1}+1}$. Then for $c\geq 2$,
    $\frac{g_{c+1}(x)}{g_{c}(x)}=2x^{2^{c-1}}\leq 2x^{2}\leq 1,$
    so $(g_{c}(x))_{c\geq 2}$ is nonincreasing. Moreover, $\frac{g_{2}(x)}{g_{1}(x)}=2x\geq 1$. Hence $g_{c}(x)\leq g_{2}(x)=4x^{3}$ for all $c\geq 1$.
    Therefore, Condition 5 implies
    $D\left(N_{c}\right)\;\geq\;D\left(M\right)+1.$
  3. Step 3:
    Monotonicity back to the protocol leaf. Because $N_{c}^{\prime}$ is a submatrix of the leaf rectangle $\widetilde{R}\times\widetilde{C}$, the Monotonicity Lemma 1 gives
    $D\left(\varepsilon\left(\widehat{M};R_{c},C_{c}\right)\right)\;\geq\;D\left(N_{c}\right)\;\geq\;D\left(M\right)+1.$
    This gives our gap.
For $c=0$, the property holds trivially.
Case c.
The row–block invariant is broken at the first step i+1i{+}1 ($B^{\prime}\leq i<B$). Let $\widetilde{M}_{i}=\varepsilon\left(\widehat{M};R_{i},C_{i}\right)$ be the unique rectangle that satisfies the invariant immediately before the (i+1)(i{+}1)-st bit is sent.
  1. Step 1:
    Row and column parameters of $\widetilde{M}_{i}$.
    By the invariant $R_{i}$ is a $(Q,T_{0})$-equipartition with
    $t=|Q|,\qquad T_{0}=\left\lceil|R|\,2^{-b}\right\rceil.$
    To break the invariant at step i+1i{+}1 the protocol must have asked at least $\left\lceil\log r\right\rceil$ row questions (each halves the number of viable blocks). Hence no more than $c\leq B-\left\lceil\log r\right\rceil$ bits have attacked the columns, so
    $\frac{|C_{i}|}{|\mathsf{Cols}(\widehat{M})|}\;\geq\;2^{-c}$
  2. Step 2:
    Relaxed $\rightarrow$ classical interlace (Lemma 10). Apply the lemma to $\widetilde{M}_{i}$ with
    $t=|Q|,\qquad x=2^{-B},\qquad y=2^{-c}.$
    $N^{\ast}\;\in\;M\left[t,2^{-B},y^{\prime}\right],\qquad y^{\prime}=2^{-c}/(1+\varepsilon).$
  3. Step 3:
    Apply condition 12.
    Invoking condition 12 with $k=i-B^{\prime}$ gives
    $D\left(N^{\ast}\right)$ $\geq D\left(M\left[t,2^{-B},2^{-c}\right]\right)$
    $\geq D\left(M\left[2,2^{-B},\left(2^{-c}\right)^{\frac{2}{t}}\right]\right)$
    $\geq D\left(M\right)-k+1.$
Implying that case c leads to complexity greater than the budget.